OpenStudy (anonymous):

a train travels 2.5 miles on a straight track with a grade of 1 degree 10'. what is the vertical rise of the train in that distance?

OpenStudy (e.mccormick):

Ah, so you have the angle of \(1^\circ20`\) 

OpenStudy (e.mccormick):

Bah, it trimmed the right... I had X over there. Then it is solve for x.

OpenStudy (anonymous):

h = 2.5 Sin[(1+1/6)degrees]=(0.0509019 *5280 ft) = 268.762 ft